OpenStudy (anonymous):

how to find the smallest positive integer which solves the following q=5(mod17) and q=8(mod12)?

OpenStudy (asnaseer):

yo can write this as follows:\[q=17m_1+5=12m_2+8\]so you need to find the smallest values for \(m_1\) and \(m_2\). we can write \(m_2\) in terms of \(m_1\) as:\[m_2=(17m_1-3)/12\]ok, so lets try increasing values of \(m_1\) starting with \(m_1=1\) and see if we can find a suitable \(m_2\):\[\begin{align}m_1&=1\implies m_2=14/12\quad&\text{reject}\\ &=2\implies m_2=31/12\quad&\text{reject}\\ &=3\implies m_2=48/12=4\quad&\text{bingo!}\\ \end{align}\] so \(m_1=3\) gives us a viable solution, which gives us:\[q=17m_1+5=17*3+5=51+5=56\]